Workman Software has 10.6 percent coupon bonds on the market with 6 years to maturity. The bonds make semiannual payments and currently sell for 109.0 percent of par.

1) What is the current yield on the bonds?

2) The YTM?

3) The effective annual yield?
Answer #1

Let the par value be $1000

Price = 109%*1000=1090

1: Current yield =Annual coupon/Price

= 10.6%*1000/1090

=0.097248

= 9.72%

2: Using financial calculator

Input:

FV=1000

PV= -1090

PMT=10.6%*1000/2= 53

N=6*2=12

Solve for I/Y as 4.323

YTM=4.3230*2= 8.65%

3: Effective annual yield = (1+YTM/n)^n-1

= (1+4.323%)^2-1

= 0.088329

=8.83%